OpenStudy (anonymous):

-7 is less than or equal to 7/3x? Solve for x please

OpenStudy (anonymous):

\[\frac{ -7 \times 3}{ 7 } \le \frac{ 7 \times 3 }{3 \times 7 } x \] \[-3 \le x\] \[x \ge - 3\]